# Scaling Laws with Vocabulary: Larger Models Deserve Larger Vocabularies

### v2

@@ 4.3 Approach 3: Parametric fit of loss formula
where $D=Hf(V)$. The first term captures the normalized loss for an ideal generative process on the data distribution. The subsequent terms reflect the effect of the non-vocabulary parameters, vocabulary parameters, and the number of training data on the loss, respectively. $E,A_{1},A_{2},B,\alpha_{1},\alpha_{2},\beta$ are learned parameters.
Fitting
We use the points ($N_{\rm nv}$, $N_{\rm v}$, $H$) collected for experiments in §4.1. Note that we do not only consider the points with the lowest loss for each FLOP budget as we want to predict loss for any combination of ($N_{\rm nv}$, $N_{\rm v}$, $H$). We add the constraint $\alpha_{1}=\beta$ following Muennighoff et al. [43]. We also filter out points with very small FLOPs following Hoffmann et al. [26]. Fitting yields $A_{1}=1.831$, $A_{2}=0.196$, $B=2.124$, $E=5.533$, $\alpha_{1}=\beta=0.447$, $\alpha_{2}=0.671$. The detailed fitting process is written in §A.5.4.
Figure 6: Vocabulary parameters of popular LLMs and predicted optimal vocabulary parameters at their reported number of training tokens, as determined by our Approach 3 (§4.3). Here we consider the practical scenarios where parameters and training data are not necessarily equally scaled. As illustrated, the vocabulary parameters remain predominantly underestimated. With the exception of Gemma2-9B, all models allocate a smaller parameter count than our predicted optimal vocabulary parameter.
Usage
After fitting the parameters in Equation 8, the optimal vocabulary size can be obtained by finding the lowest loss w.r.t the vocabulary size, with a constraint of FLOPs budget. For example, given $N_{\rm nv}$ and FLOPs budget $C$ , by replacing $[Hf(V)]$ with $C/(6(N_{\rm nv}+N_{v}))$ and finding the solution of $\frac{\partial\mathcal{L}_{u}}{\partial V}=0$ via numerical search, we can get the prediction. The details of $\frac{\partial\mathcal{L}_{u}}{\partial V}$ is written in §A.2. Note that all of the proposed approaches can be used in optimally allocating ($N_{\rm nv},N_{\rm v},H$) altogether, while Approach 3 is more flexible in predicting the locally optimal $N_{\rm v}$ when ($N_{\rm nv}$, $H$) are not following the Chinchilla’s law [26], i.e. equally-scaled law. The reason is that the loss formula in Approach 3 does not only considers the combinations ($N_{\rm nv},N_{\rm v},H$) which reach the optimal given a certain training budget. By fixing $N_{\rm nv}$ and varying $C$ in Approach 3, we can predict the locally optimal vocabulary size with different amount of training characters. This property makes Approach 3 more valuable, since modern LLMs [67, 64, 3, 4, 7] usually leverage overly sufficient training data to build powerful models with relatively low inference costs.
@@ 5 Discussion
Table 1 reports the predicted optimal vocabulary parameters and sizes based on the proposed three approaches, where the amount of training data is optimally allocated, i.e. equally scaled with the non-vocabulary parameters [26]. As shown in Figure 1, the predictions from all proposed approaches align closely. $N_{\rm nv}$ should be scaled faster than $N_{\rm v}$. Notably, mainstream LLMs typically assign fewer parameters to vocabulary than what is optimal. However, the community is starting to shift to larger vocabularies, such as with Llama3 [40] having a 128K vocabulary size up from 32K of Llama2 [67]. However, scaling data is still the most critical part, and solving data scarcity issues should be a focus of future work [69].
@@ 6 Related work
Large language models
The Transformer [68] has proven to be a very scalable architecture with consistent performance gains which has led to a series of large language models (LLMs) [11, 14, 51, 46, 20, 29, 52, 67, 70, 40, 8, 4, 37, 25, 61, 64, 7, 38, 32, 83]. Through their training to predict subsequent tokens in a sequence, these models acquire a deep understanding of language enabling them to perform a variety of language tasks directly after pre-training. Their capabilities include code generation [33, 3, 42, 82, 81], mathematical reasoning [72, 5], question answering [47, 44] among others. In our work, we pre-train large language models from scratch on English corpora and focus on their loss during training and downstream performance on common tasks after training.
@@
Scaling laws aim to develop a predictive framework to find the best allocation of compute resources to maximize model performance. Besides language models, they have been studied for diffusion models [39], visual auto-regressive modeling [65] and contrastive language-image learning [13]. For language models, Kaplan et al. [30] show that model performance improves as a power law with more compute allocated to both parameters or data. Hoffmann et al. [26] show that the allocation of compute should be such that parameters and data are scaled equally. Other work investigates predicting other attributes such as downstream performance [23, 28, 54] or considering inference time [56]. Some research also predicts the expected benefit and optimal allocation under data constraints [43, 75]. Across all these works, the vocabulary size has generally been ignored. Kaplan et al. [30] even explicitly only consider non-embedding parameters. Our work highlights the critical role of vocabulary in scaling LLMs. Through our predictive frameworks for determining the optimal vocabulary size, we hope that future research will pay more attention to the vocabulary.
Vocabulary in language models
The vocabulary of a language model influences its performance significantly [62, 71, 74]. A larger vocabulary size helps cover more words thus reducing the likelihood of out-of-vocabulary (OOV) cases [21]. Takahashi and Tanaka-Ishii [62] find that larger vocabularies are better at capturing the true statistical distribution of language. Similarly, expanding vocabulary in multilingual models [71, 15, 80, 34] improves performance, especially for low-resource languages. However, large vocabularies [31] increase the computational overhead during both training and generation phases. For example, Liao et al. [35] demonstrate that low-frequency words, due to their limited occurrence in training data, often lack sufficient examples to develop robust representations when vocabularies are excessively large. Dou et al. [19] reveal that expanding vocabularies during continual pre-training can lead to significant performance degradation for low-resource languages. More recently, Dagan et al. [18] explored the trade-offs associated with vocabulary size, proposing optimal vocabulary sizes for both memory efficiency and inference speed in code generation tasks. Our work complements these efforts by focusing on the broader impact of vocabulary size on downstream performance across various tasks. Specifically, we address a critical, under-explored question: How can we optimally allocate vocabulary size to maximize the downstream performance with the same compute budget?
Byte-level language models
@@ 7 Conclusion
We investigate the impact of the vocabulary size when scaling language models. We analyze and verify that there exists an optimal vocabulary size for a given FLOPs budget. Subsequently, we develop 3 approaches to predict the optimal vocabulary size. Our first approach uses a set of empirical training runs across different IsoFLOPs regimes to fit a scaling law. The second approach investigates the FLOPs w.r.t. the vocabulary size and estimates the vocabulary size with derivatives. The third approach consists of a parametric function to directly predict the impact of different attributes on loss. Across all approaches, we find that while vocabulary parameters should be scaled slower than other parameters, they are still critical for performance and we can accurately predict their optimal allocation. We make predictions for larger models and empirically verify our approaches on up to 3B parameters and on varying amounts of training data. Our results show that models trained with an optimal vocabulary size as predicted by our approaches outperform models with a conventional vocabulary size under the same FLOPs budget.
@@ Appendix A Appendix
$\frac{\partial C}{\partial V}$ $=\frac{\partial}{\partial V}\left[6(N_{nv}+dV)H\left(f(V)\right)\right]$
$=\frac{\partial}{\partial V}\left[6(N_{nv}+dV)H\left(a(\log(V))^{2}+b\log(V)+c\right)\right]$
$=6H\Bigg{[}(N_{nv}+dV)\frac{d}{dV}\left(a(\log(V))^{2}+b\log(V)+c\right)$
$\quad\quad\quad\quad+\left(a(\log(V))^{2}+b\log(V)+c\right)\frac{d}{dV}(N_{nv}+dV)\Bigg{]}$
$=6H\Bigg{[}(N_{nv}+Vd)\frac{2a\log(V)+b}{V}+\left(a(\log(V))^{2}+b\log(V)+c\right)d\Bigg{]}.$
@@
Figure 8: Left: The curve of the derivative of FLOPs with respect to vocabulary size $V$. The curve of $\frac{\partial C}{\partial V}$ increases as $V$ increases, and the FLOPs reach a minima at the solution of $\frac{\partial C}{\partial V}=0$. Middle: The curve of FLOPs with respect to vocabulary size $V$, where $V$ reaches its optimal point $V$. Right: The curve of training characters with a given FLOPs budget. Take $N_{nv}=302M$ and $H=43B$ as an example. The FLOPs budget is decided by the $N_{nv}$, $H$ and the predicted $V$.
@@
Here we provide how we derive the loss w.r.t the vocabulary size given a FLOPs budget $C$ in Approach 3. After substituting the $[Hf(V)]$ with the $C/(6(N_{nv}+N_{v})$ based on Equation 9:
$\mathcal{L}_{u}=-E+\frac{A_{1}}{N_{nv}^{\alpha_{1}}}+\frac{A_{2}}{N_{v}^{\alpha_{2}}}+\frac{B}{[C/(6(N_{nv}+N_{v})]^{\beta}}.$ (10)
The loss is solely dependent on the $N_{v}=Vd$, given a $N_{nv}$. The derivative w.r.t. $V$ is:
$\frac{\partial\mathcal{L}_{u}}{\partial V}$ $=\frac{\partial}{\partial V}\left(\frac{A_{2}}{(Vd)^{\alpha_{2}}}\right)+\frac{\partial}{\partial V}\left(\frac{B}{\left(\frac{C}{6(N_{nv}+Vd)}\right)^{\beta}}\right)$
$=-\alpha_{2}\frac{A_{2}d}{(Vd)^{\alpha_{2}+1}}+\beta\frac{B\frac{Cd}{6(N_{nv}+Vd)^{2}}}{\left(\frac{F}{6(N_{nv}+Vd)}\right)^{\beta+1}}.$
@@ A.4 Exploration of Larger Range of Vocabulary Sizes
Figure 10: Loss curves with larger range of vocabulary sizes (from [4K, 96K] to [1K, 512K]), given a certain FLOPs budget. The model performance degrades consistently when the vocabulary size goes beyond the optimal configuration.
@@
To further verify that there is always an optimal vocabulary size holds for a larger range of vocabulary lists, we increase the range of vocabulary sizes from 1K to 512K, with the $N_{nv}$ fixed as 33M. As depicted in the Figure 10, the model’s performance declines consistently as the vocabulary size increases beyond the optimal configuration. This figure shows loss curves for vocabulary sizes up to 512K, given a specific FLOPs budget. The data indicates a consistent degradation in model performance with the vocabulary size away from the optimal one. It suggests that there is a critical point beyond which the model’s efficiency in handling the vocabulary diminishes. This exploration underscores the importance of carefully selecting the vocabulary size to maintain optimal model performance within the constraints of a given computational budget.
A.5 Implementation details
A.5.1 Setting of model architecture, vocabulary size and training characters
We list the architectures of the models and the corresponding number of training characters in Table 4. For each model family, we fix the non-vocabulary parameters $N_{nv}$ and vary the vocabulary size. We adopt the Llama architecture [66], except for the vocabulary size. For the vocabulary size, we use numbers divisible by 128 for compatibility with NVIDIA’s tensor core to accelerate matrix multiplication  [^3] https://docs.nvidia.com/deeplearning/performance/dl-performance-matrix-multiplication/index.html. Specifically, the vocabulary sizes we adopt for each model family are 4096, 6144, 8192, 10240, 16384, 24576, 32768, 48128, 64512 and 96256. The expected number of training tokens $D$ and characters $H$ vary slightly given a fixed number of non-vocabulary parameters and a FLOP budget. We use the middle-sized $V$ of 16384 to determine the number of training characters and the corresponding FLOPs budget, except for $N_{nv}=2870M$ we use $V=32K$.
@@ A.5.1 Setting of model architecture, vocabulary size and training characters
$N_{nv}$ (M) #Sequence Length #Layers #Heads #Embedding Dim. #Intermediate Size Training Characters (B)
@@
A.5.2 The relationship between non-vocabulary parameters and embedding dimension
According to the observation in Kaplan et al. [30], the depth-width ratio has a relatively small effect on performance given the total non-vocabulary parameters. Thus, to ease the modeling of our scaling laws taking vocabulary size into account, we take the width (i.e. embedding dimension) as given following prior work [30, 26, 43, 67, 79]. The relationship between the non-vocabulary parameters $N_{nv}$ and embedding dimension $d$ used in our experiments are in Table 5.
Table 5: The relationship between the non-vocabulary parameters $N_{nv}$ and the embedding dimension used in our experiments.
Non-vocabulary Parameters $N_{nv}$Embedding Dim.
$N_{nv}$ $\leq 50M$ 512
$50M<N_{nv}\leq 200M$ 768
$200M<N_{nv}\leq 500M$ 1024
$500M<N_{nv}\leq 1B$ 1536
$1B<N_{nv}\leq 2B$ 2048
$2B<N_{nv}\leq 5B$ 3200
$5B<N_{nv}\leq 10B$ 4096
$10B<N_{nv}\leq 20B$ 5120
$20B<N_{nv}\leq 50B$ 6048
$50B<N_{nv}\leq 100B$ 8192
$100B<N_{nv}\leq 200B$ 12288
$200B<N_{nv}\leq 500B$ 16384
$500B<N_{nv}\leq 1000B$ 20480
A.5.3 Training details
The maximum learning rate is set to 4e-4 and decays to 10% i.e. 4e-5 similar to prior scaling work [26, 43]. We use AdamW [36] as our optimizer and accelerate training with bfloat16 mixed precision training. For models with $N_{nv}<1130M$, we use a single node with 8 GPUs for training. Otherwise, we adopt the Megatron-LM framework [59] for multi-node training with 8 GPUs on each node. For our experiments with $N_{nv}=2870M$, it takes about 120 hours to train on over $500B$ training characters with 64 total GPUs. We use a global batch size of 512 for all runs and run all experiments on 40GB Nvidia-A100 GPUs.
A.5.4 Fitting techniques

### v3

@@ 4.3 Approach 3: Parametric fit of loss formula
where $D=Hf(V)$. The first term captures the normalized loss for an ideal generative process. The subsequent terms reflect the effect of the non-vocabulary parameters, vocabulary parameters, and the number of training data on the loss, respectively. $E,A_{1},A_{2},B,\alpha_{1},\alpha_{2},\beta$ are learned parameters.
Fitting    We use the points ($N_{\rm nv}$, $N_{\rm v}$, $H$) collected for experiments in §4.1. Note that we do not only consider the points with the lowest loss for each FLOP budget as we want to predict loss for any combination of ($N_{\rm nv}$, $N_{\rm v}$, $H$). We add the constraint $\alpha_{1}=\beta$ following Muennighoff et al. [44]. We also filter out points with very small FLOPs following Hoffmann et al. [26]. Fitting yields $A_{1}=1.831$, $A_{2}=0.196$, $B=2.124$, $E=5.533$, $\alpha_{1}=\beta=0.447$, $\alpha_{2}=0.671$. The detailed fitting process is written in §A.7.4.
Usage    After fitting the parameters in Equation 8, the optimal vocabulary size can be obtained by finding the lowest loss w.r.t the vocabulary size, with a constraint of FLOPs budget. For example, given $N_{\rm nv}$ and FLOPs budget $C$ , by replacing $[Hf(V)]$ with $C/(6(N_{\rm nv}+N_{v}))$ and finding the solution of $\frac{\partial\mathcal{L}_{u}}{\partial V}=0$ via numerical search, we can get the prediction. The details of $\frac{\partial\mathcal{L}_{u}}{\partial V}$ is written in §A.2. Note that all of the proposed approaches can be used in optimally allocating ($N_{\rm nv},N_{\rm v},H$) altogether, while Approach 3 is more flexible in predicting the locally optimal $N_{\rm v}$ when ($N_{\rm nv}$, $H$) are not following the Chinchilla’s law [26], i.e. equally-scaled law. The reason is that the loss formula in Approach 3 does not only considers the combinations ($N_{\rm nv},N_{\rm v},H$) which reach the optimal given a certain training budget. By fixing $N_{\rm nv}$ and varying $C$ in Approach 3, we can predict the locally optimal vocabulary size with different amount of training characters. This property makes Approach 3 more valuable, since modern LLMs [70, 67, 3, 4, 7] usually leverage overly sufficient training data to build powerful models with relatively low inference costs.
Figure 6: Vocabulary parameters of popular LLMs and predicted optimal vocabulary parameters at their reported number of training tokens, as determined by our Approach 3 (§4.3). Here we consider the practical scenarios where parameters and training data are not necessarily equally scaled. As illustrated, the vocabulary parameters remain predominantly underestimated. With the exception of Gemma2-9B, all models allocate a smaller vocabulary parameter count than our prediction.
@@ 5 Discussion
Table 1 reports the predicted optimal vocabulary parameters and sizes based on the proposed three approaches, where the amount of training data is optimally allocated, i.e. equally scaled with the non-vocabulary parameters [26]. Aligned with the trend shown in Figure 1, the predictions from all proposed approaches align closely. $N_{\rm nv}$ should be scaled faster than $N_{\rm v}$. Notably, mainstream LLMs typically assign fewer parameters to vocabulary than what is optimal. However, the community is starting to shift to larger vocabularies, such as with Llama3 [41] having a 128K vocabulary size up from 32K of Llama2 [70]. However, scaling data is still the most critical part, and solving data scarcity issues should be a focus of future work [72].
@@ 6 Related work
Language models
The Transformer [71] has proven to be a scalable architecture for language models, especially large language model (LLMs) [11, 14, 52, 47, 20, 29, 53, 70, 73, 41, 8, 4, 38, 25, 62, 67, 7, 39, 32, 88]. These models typically acquire a deep understanding of language enabling them to perform multiple tasks after a pre-training period and an optional fine-tuning period. Their capabilities include code generation [33, 3, 43, 87, 86], mathematical reasoning [76, 5], question answering [48, 45] among others. Given the expensive deployment costs required by the language models, various techniques can be adopted for efficient inference [64, 65, 78, 74, 36]. In our work, we pre-train large language models from scratch on English corpora and focus on their validation loss and downstream performance after training.
@@
Scaling laws aim to develop a predictive framework to find the best allocation of compute resources to maximize model performance. Besides language models, they have been studied in other domains[40, 68, 13]. For language models, Kaplan et al. [30] show that performance improves as a power law with more compute allocated to both parameters or data. Hoffmann et al. [26] show that the compute allocation of parameters and data should be scaled equally. Other work considers various cases such as downstream performance [23, 28, 55], inference time [57] or data constraints [44, 80]. However, the effect of vocabulary size has generally been ignored previously.
@@ 7 Conclusion
We investigate the impact of the vocabulary size in language models. We analyze and verify that there exists an optimal vocabulary size for a given FLOPs budget. Subsequently, we develop 3 approaches to predict the optimal vocabulary size. Our first approach uses empirical training runs across different IsoFLOPs regimes to fit a scaling law. The second approach investigates the FLOPs w.r.t. the vocabulary size and estimates the vocabulary size with derivatives. The third approach consists of a parametric function to predict the impact of different attributes on loss. Across all approaches, we find that while vocabulary parameters should be scaled slower than other parameters, they are still critical for performance and we can accurately predict their optimal allocation. We make predictions for larger models and empirically verify our approaches on up to 3B parameters and on varying amounts of training data. We show that models trained with an optimal vocabulary size as predicted by our approaches outperform models with a conventional vocabulary size under the same FLOPs budget.
Acknowledgements
This work was supported by in part by the Theme-based Research Scheme (TRS) project T45-701/22-R of the Research Grants Council (RGC), Hong Kong SAR.
@@ Appendix A Appendix
$\frac{\partial C}{\partial V}$ $=\frac{\partial}{\partial V}\left[6(N_{\rm nv}+dV)H\left(f(V)\right)\right]$
$=\frac{\partial}{\partial V}\left[6(N_{\rm nv}+dV)H\left(a(\log(V))^{2}+b\log(V)+c\right)\right]$
$=6H\Bigg{[}(N_{\rm nv}+dV)\frac{d}{dV}\left(a(\log(V))^{2}+b\log(V)+c\right)$
$\quad\quad\quad\quad+\left(a(\log(V))^{2}+b\log(V)+c\right)\frac{d}{dV}(N_{\rm nv}+dV)\Bigg{]}$
$=6H\Bigg{[}(N_{\rm nv}+Vd)\frac{2a\log(V)+b}{V}+\left(a(\log(V))^{2}+b\log(V)+c\right)d\Bigg{]}.$
@@
Figure 8: Left: The curve of the derivative of FLOPs with respect to vocabulary size $V$. The curve of $\frac{\partial C}{\partial V}$ increases as $V$ increases, and the FLOPs reach a minima at the solution of $\frac{\partial C}{\partial V}=0$. Middle: The curve of FLOPs with respect to vocabulary size $V$, where $V$ reaches its optimal point $V$. Right: The curve of training characters with a given FLOPs budget. Take $N_{\rm nv}=302M$ and $H=43B$ as an example. The FLOPs budget is decided by the $N_{\rm nv}$, $H$ and the predicted $V$.
@@
Here we provide how we derive the loss w.r.t the vocabulary size given a FLOPs budget $C$ in Approach 3. After substituting the $[Hf(V)]$ with the $C/(6(N_{\rm nv}+N_{v})$ based on Equation 9:
$\mathcal{L}_{u}=-E+\frac{A_{1}}{N_{\rm nv}^{\alpha_{1}}}+\frac{A_{2}}{N_{v}^{\alpha_{2}}}+\frac{B}{[C/(6(N_{\rm nv}+N_{v})]^{\beta}}.$ (10)
The loss is solely dependent on the $N_{v}=Vd$, given a $N_{\rm nv}$. The derivative w.r.t. $V$ is:
$\frac{\partial\mathcal{L}_{u}}{\partial V}$ $=\frac{\partial}{\partial V}\left(\frac{A_{2}}{(Vd)^{\alpha_{2}}}\right)+\frac{\partial}{\partial V}\left(\frac{B}{\left(\frac{C}{6(N_{\rm nv}+Vd)}\right)^{\beta}}\right)$
$=-\alpha_{2}\frac{A_{2}d}{(Vd)^{\alpha_{2}+1}}+\beta\frac{B\frac{Cd}{6(N_{\rm nv}+Vd)^{2}}}{\left(\frac{F}{6(N_{\rm nv}+Vd)}\right)^{\beta+1}}.$
@@ A.4 Exploration of Larger Range of Vocabulary Sizes
Figure 10: Loss curves with larger range of vocabulary sizes (from [4K, 96K] to [0.5K, 512K]), given a certain FLOPs budget. The model performance degrades consistently when the vocabulary size goes beyond the optimal configuration.
@@
To further verify that there is always an optimal vocabulary size holds for a larger range of vocabulary lists, we increase the range of vocabulary sizes from 0.5K to 512K, with the $N_{\rm nv}$ fixed as 33M. As depicted in the Figure 10, the model’s performance declines consistently as the vocabulary size increases beyond the optimal configuration. This figure shows loss curves for vocabulary sizes up to 512K, given a specific FLOPs budget. The data indicates a consistent degradation in model performance with the vocabulary size away from the optimal one. It suggests that there is a critical point beyond which the model’s efficiency in handling the vocabulary diminishes. This exploration underscores the importance of carefully selecting the vocabulary size to maintain optimal model performance within the constraints of a given computational budget.
A.5 The Vocabulary-insensitive Metric: $\mathcal{L}_{u}$ and BPC
Figure 11: Correlation between the unigram-normalized loss $\mathcal{L}_{u}$ and BPC, where $\rho$ and $e$ denote the Pearson correlation coefficient and the root mean square error of the linear fit, respectively.
BPC reflects the ability to compress external text corpora [27], while the unigram-normalized loss reflects the model’s ability to predict tokens normalized by the token frequency. Figure 11 shows the relationship between the Unigram-normalized Loss $\mathcal{L}_{u}$ and Bits Per Character (BPC) with a linear fit. We select the models of the final training steps for each $N_{nv}$ and each vocabulary size. The high correlation coefficient ($\rho$ = 0.9888) and low error ($e$ = 0.0683) indicate a strong linear relationship between these two metrics generally. However, it exists slight different trends due to different normalizations.
A.6 More Explanations about Why We Separate Vocabulary Parameters from the total Model Parameters
Non-vocabulary parameters can benefit from increases in both depth and width, allowing for more complex hierarchical representations and broader feature capture. In contrast, vocabulary parameters, associated with word embeddings and language model heads, are generally confined to a single layer, limiting their ability to benefit from increases in the model depth. Their primary avenue for expansion is through increasing the width. This disparity in growth potential between non-vocabulary and vocabulary parameters suggests that to maintain a balanced growth rate, it may be necessary to expand the vocabulary size along with the depth. This would allow the vocabulary parameters to keep pace with the growth of non-vocabulary parameters.
A.7 Implementation details
A.7.1 Setting of model architecture, vocabulary size and training characters
We list the architectures of the models and the corresponding number of training characters in Table 4. For each model family, we fix the non-vocabulary parameters $N_{\rm nv}$ and vary the vocabulary size. We adopt the Llama architecture [69], except for the vocabulary size. For the vocabulary size, we use numbers divisible by 128 for compatibility with NVIDIA’s tensor core to accelerate matrix multiplication  [^3] https://docs.nvidia.com/deeplearning/performance/dl-performance-matrix-multiplication/index.html. Specifically, the vocabulary sizes we adopt for each model family are 4096, 6144, 8192, 10240, 16384, 24576, 32768, 48128, 64512 and 96256. The expected number of training tokens $D$ and characters $H$ vary slightly given a fixed number of non-vocabulary parameters and a FLOP budget. We use the middle-sized $V$ of 16384 to determine the number of training characters and the corresponding FLOPs budget, except for $N_{\rm nv}=2870M$ we use $V=32K$.
@@ A.7.1 Setting of model architecture, vocabulary size and training characters
$N_{\rm nv}$ (M) #Sequence Length #Layers #Heads #Embedding Dim. #Intermediate Size Training Characters (B)
@@
A.7.2 The relationship between non-vocabulary parameters and embedding dimension
According to the observation in Kaplan et al. [30], the depth-width ratio has a relatively small effect on performance given the total non-vocabulary parameters. Thus, to ease the modeling of our scaling laws taking vocabulary size into account, we take the width (i.e. embedding dimension) as given following prior work [30, 26, 44, 70, 84]. The relationship between the non-vocabulary parameters $N_{\rm nv}$ and embedding dimension $d$ used in our experiments are in Table 5.
Table 5: The relationship between the non-vocabulary parameters $N_{\rm nv}$ and the embedding dimension used in our experiments.
Non-vocabulary Parameters $N_{\rm nv}$Embedding Dim.
$N_{\rm nv}$ $\leq 50M$ 512
$50M<N_{\rm nv}\leq 200M$ 768
$200M<N_{\rm nv}\leq 500M$ 1024
$500M<N_{\rm nv}\leq 1B$ 1536
$1B<N_{\rm nv}\leq 2B$ 2048
$2B<N_{\rm nv}\leq 5B$ 3200
$5B<N_{\rm nv}\leq 10B$ 4096
$10B<N_{\rm nv}\leq 20B$ 5120
$20B<N_{\rm nv}\leq 50B$ 6048
$50B<N_{\rm nv}\leq 100B$ 8192
$100B<N_{\rm nv}\leq 200B$ 12288
$200B<N_{\rm nv}\leq 500B$ 16384
$500B<N_{\rm nv}\leq 1000B$ 20480
A.7.3 Training details
The maximum learning rate is set to 4e-4 and decays to 10% i.e. 4e-5 similar to prior scaling work [26, 44]. We use AdamW [37] as our optimizer and accelerate training with bfloat16 mixed precision training. For models with $N_{\rm nv}<1130M$, we use a single node with 8 GPUs for training. Otherwise, we adopt the Megatron-LM framework [60] for multi-node training with 8 GPUs on each node. For our experiments with $N_{\rm nv}=2870M$, it takes about 120 hours to train on over $500B$ training characters with 64 total GPUs. We use a global batch size of 512 for all runs and run all experiments on 40GB Nvidia-A100 GPUs.
A.7.4 Fitting techniques